ou have observed the following returns over time:

Year Stock X Stock Y Market
2014 13 % 13 % 14 %
2015 20 5 9
2016 -16 -2 -12
2017 3 3 3
2018 24 12 16

Assume that the risk-free rate is 7% and the market risk premium is 5%.

  1. What are the betas of Stocks X and Y? Do not round intermediate calculations. Round your answers to two decimal places.

    Stock X:

    Stock Y:

  2. What are the required rates of return on Stocks X and Y? Do not round intermediate calculations. Round your answers to two decimal places.

    Stock X:   %

    Stock Y:   %

  3. What is the required rate of return on a portfolio consisting of 80% of Stock X and 20% of Stock Y? Do not round intermediate calculations. Round your answer to two decimal places.

      %
Answer #1
Given,
Year Return on stock X (Rx) Return on stock Y (Ry) Return on market (Rm)
2014 13% 13% 14%
2015 20% 5% 9%
2016 -16% -2% -12%
2017 3% 3% 3%
2018 24% 12% 16%
Rf=7%
Market Risk premium=5%
a) Calculation of Betas
We know beta of stock= (Expected return on stock-Rf)/(Expected return from market-Rf)
Betas of stock X
2014 =(13-7)/(14-7) 0.86
2015 =(20-7)/(9-7) 6.50
2016 =(-16-7)/(-12-7) 1.21
2017 =(3-7)/(3-7) 1.00
2018 =(24-7)/(16-7) 1.89
Betas of stock Y
2014 =(13-7)/(14-7) 0.86
2015 =(5-7)/(9-7) -1.00
2016 =(-2-7)/(-12-7) 0.47
2017 =(3-7)/(3-7) 1.00
2018 =(12-7)/(16-7) 0.56
b) Calculation of rate of return on stocks
As per CAPM,
Required return= Rf+Market risk premium*Beta
Rate of return on stock X
2014 =7+5*0.86 11.30%
2015 =7+5*6.50 39.50%
2016 =7+5*1.21 13.05%
2017 =7+5*1 12.00%
2018 =7+5*1.89 16.45%
Rate of return on stock Y
2014 =7+5*0.86 11.30%
2015 =7+5*-1 2.00%
2016 =7+5*0.47 9.35%
2017 =7+5*1 12.00%
2018 =7+5*0.56 9.80%
c) Required return on portfolio
Year Return on stock X (Rx) Weight of X (Wx) Rx*Wx Return on stock Y (Ry) Weight of X (Wy) Ry*Wy Expected return on portfolio
2014 13% 0.80 10.40% 13% 0.20 2.60% =10.40+2.60 13.00%
2015 20% 0.80 16.00% 5% 0.20 1.00% =16+1 17.00%
2016 -16% 0.80 -12.80% -2% 0.20 -0.40% -12.80-0.40 -13.20%
2017 3% 0.80 2.40% 3% 0.20 0.60% =2.40+0.60 3.00%
2018 24% 0.80 19.20% 12% 0.20 2.40% =19.20+2.40 21.60%
Expected return of portfolio= Weighted average (See table above)